OpenStudy (anonymous):

Simplify: (sin Θ − cos Θ)2 - (sin Θ - cos Θ)2

OpenStudy (anonymous):

0 2 sin^2sigma Cos^2sigma

OpenStudy (anonymous):

@alyssa_xo

OpenStudy (anonymous):

oh and those twos on the () are ^2

OpenStudy (anonymous):

I tried to do it the way you showed me but I kept doing something wrong

OpenStudy (alyssa_xo):

you're in bio btw, you can switch subjects by hovering over http://puu.sh/i1Vm4/6075c5fc3e.jpg and either clicking your new subject or clicking find more subjects

OpenStudy (anonymous):

oh shoot

OpenStudy (alyssa_xo):

nobody cares lol

OpenStudy (alyssa_xo):

just a heads up when you ask your next question so you can get the appropriate help

OpenStudy (anonymous):

haha ill remember that!

OpenStudy (alyssa_xo):

ok so let's see \((a-b)^2=a^2-2ab+b^2\\(a-b)^2-(a-b)^2=(a^2-2ab+b^2)-(a^2-2ab+b^2)\) this is probably where you went wrong, did you distribute the minus sign correctly?

OpenStudy (anonymous):

nope it doesnt look like it haha

OpenStudy (alyssa_xo):

note \((a^2-2ab+b^2)-(a^2-2ab+b^2)=\\(a^2-2ab+b^2)-a^2-(-2ab)-b^2\)

OpenStudy (anonymous):

okay, that makes sense

OpenStudy (alyssa_xo):

and so we have \(a^2-2ab+b^2-a^2+2ab-b^2\)

OpenStudy (alyssa_xo):

looks like 0 to me

OpenStudy (alyssa_xo):

wow, lmao a 6 year old could have done this

OpenStudy (anonymous):

Oh wow thanks for your help! And for being mean lol

OpenStudy (alyssa_xo):

lol didn't mean to be mean, just to point out that we did unneeded work

OpenStudy (alyssa_xo):

look at the question, it's of the form \(a^2-a^2\) anything minus itself = 0

OpenStudy (alyssa_xo):

so we didn't have to expand the stuff

OpenStudy (anonymous):

ahhhhh good point lol

OpenStudy (anonymous):

thanks!